An acidic solution is 2 mM in each of the following metal ions: Zn2+, Cu2+, Co2+, Ca2+.

a) Which of the metals would precipitate as their hydroxides at pH = 6.00? (Select all that apply. See theappendix.)

-zinc

-copper

-cobalt

-calcium

b) Which of the metals would precipitate as their hydroxides at pH = 8.00?

-zinc

-copper

-cobalt

-calcium

c) At what pH do the metal hydroxides begin to participate?

Zn(OH)2 ____

Cu(OH)2 ____

Co(OH)2 ____

Ca(OH)2 ____

Substance Ksp at 25 oC
Aluminum
Al(OH)3 1.9 x 10-33
Barium
Ba(OH)2 5.0 x 10-3
BaCO3 8.1 x 10-9
BaSO4 1.1 x 10-10
Ba3(PO4)2 3.4 x 10-23
Cadmium
Cd(OH)2 2.5 x 10-14
CdCO3 5.2 x 10-12
CdS 8.0 x 10-27
Calcium
Ca(OH)2 5.5 x 10-6
CaCO3 4.8 x 10-9
Ca3(PO4)2 1.0 x 10-26
CaF2 3.9 x 10-11
Chromium
Cr(OH)3 6.3 x 10-31
Cobalt
Co(OH)2 1.6 x 10-15
CoS 4.0 x 10-21
Copper
Cu(OH)2 2.2 x 10-20
Cu2S 1.6 x 10-48
CuCO3 1.4 x 10-10
CuS 6.3 x 10-36
Iron
Fe(OH)2 8.0 x 10-16
FeS 6.3 x 10-18
Fe(OH)3 2.5 x 10-39
FePO4 9.9 x 10-29
Lead
Pb(OH)2 2.8 x 10-16
PbF2 3.7 x 10-8
PbCl2 1.7 x 10-5
PbBr2 6.3 x 10-6
PbI2 6.5 x 10-9
PbCrO4 1.8 x 10-14
PbSO4 1.7 x 10-8
PbS 8.4 x 10-28
Magnesium
Mg(OH)2 1.8 x 10-11
Manganese
Mn(OH)2 1.9 x 10-13
MnCO3 8.8 x 10-11
MnS 5.6 x 10-16
Nickel
Ni(OH)2 1.6 x 10-14
Silver
AgCl 1.8 x 10-10
AgBr 5.0 x 10-13
AgI 8.3 x 10-17
AgCN 1.2 x 10-16
Ag2S 6.3 x 10-50
Ag2CrO4 1.1 x 10-12
Ag3PO4 2.6 x 10-18
Tin
Sn(OH)2 1.4 x 10-28
SnS 1.3 x 10-23
Zinc
ZnCO3 1.0 x 10-10
Zn(OH)2 4.5 x 10-17
ZnS 1.1 x 10-21
Answer #1

1) We can determine that ppt will form or not by taking into account the ksp

-zinc

Zn(OH)2 -----> Zn2+ + 2OH-
Ksp = [Zn2+][OH-]^2

To form a precipitate, the ionic product must be greater than the Ksp.

4.5 x 10-17 = [Zn+2][OH-]^2

Zn +2 = 1/2 OH-
1/2[OH-]^3 = = 4.5 x 10-17
[OH] = 11.61 x 10^-6M
pOH = - log (6.7 x 10^-6) =4.93
pH = 14.00 - pOH = 14.00 - 4.93 = 9.06

So minimum pH required is 8.82 for Zn+2 so it will not ppt out

-copper

Cu(OH)2 -----> Cu2+ + 2OH-
Ksp = [Cu2+][OH-]^2

To form a precipitate, the ionic product must be greater than the Ksp.

2.2 x 10-20= [Cu+2][OH-]^2

Cu +2 = 1/2 OH-
1/2[OH-]^3 = 2.2 x 10-20

[OH] = 4.69 x 10^-7M
pOH = - log (4.69 x 10^-7) = 6.32
pH = 14.00 - pOH = 14.00 - 4.72 = 7.67

So minimum pH required is 7.67 for Cu+2 so it will not ppt out

-cobalt

Co(OH)2 -----> Co2+ + 2OH-
Ksp = [Co2+][OH-]^2

To form a precipitate, the ionic product must be greater than the Ksp.

1.6 x 10-15= [Co+2][OH-]^2

Co +2 = 1/2 OH-
1/2[OH-]^3 = 1.6 x 10-15

[OH] = 2.4 x 10^-5M
pOH = - log (2.4 x 10^-5) = 4.61

pH = 14.00 - pOH = 14.00 - 4.61= 9.38

So minimum pH required is 9.38 for Co+2 so it will not ppt out

-calcium

Ca(OH)2 -----> Ca2+ + 2OH-
Ksp = [Ca2+][OH-]^2

To form a precipitate, the ionic product must be greater than the Ksp.

5.5 x 10-6 = [Ca+2][OH-]^2

Ca +2 = 1/2 OH-
1/2[OH-]^3 = 5.5 x 10-6

[OH] = 3.31 x 10^-2M
pOH = - log (3.31 x 10^-2) = 1.48

pH = 14.00 - pOH = 14.00